Which of the following statements is/are correct? I. At 6% interest, the present value of: $400 for the first year, $600 for the second year, and $800 for the third year is $1,603.00. II. The future value of the following mixed cash flow stream (if it is from an annuity due at 6% interest):$400 for the first year, $600 for the second year, and $800 for the third year is $1,999 (rounded).
Answer #1

Anwer to option1

Statement is wrong correct present value is 1570

Present value formula is =C/(1+i)^n

C=cash flow

I=rate of interest (discounting rate)

n=no of discounting periods

Caluclotion

400/1.06+600/(1.06)^2+ 800/(1.06)^3=1583

Explanation to answer 2

Answer is correct

Calculation 400*(1.06)^3+600*(1.06)^2+800*(1.06)=$1999